# From No-Regret to Strategically Robust Learning in Repeated Auctions

### v1

@@ 4.2 Main lemma
We introduce some notation to simplify the lemma statement. We let $\mathbf{e}:=(1,0,\dots,0)$ denote the first standard basis vector in $\mathbb{R}^{K+1}$. For each round $t\in[T]$, we let $Q^{(t)}:=(\bm{\pi}^{(1,t)},\dots,\bm{\pi}^{(n,t)})$ be the profile of bidders’ quantile strategies, which determine their bidding strategies $s_{i}^{(t)}$, and we let $\textnormal{Rev}_{\mathcal{M}^{(t)}}\big(\mathcal{D}\mid Q^{(t)}\big)$ denote the auctioneer’s expected revenue at round $t$ conditioned on $Q^{(t)}$, i.e.,
@@
Now we state the main lemma (Lemma 4.7), which relates the gradient of the quantile utility function to the difference between the auctioneer’s expected revenue at round $t$ and the expected revenue of the auxiliary auction. The special case of this relation for single-bidder first-price auctions was previously shown by Kumar et al. (2024, page 21).
@@
We clarify that in Algorithm 1, bidder $i$’s quantile utility function $q_{i}^{(t)}$ at round $t$ (as defined in Eq. (7)) depends on the other bidders’ bids $\mathbf{b}_{-i}^{(t)}$, which in turn are determined by the other bidders’ quantile/bidding strategies and private values $\mathbf{v}_{-i}^{(t)}$ at round $t$. Hence, the expectation in Eq. (15) is taken with respect to the randomness of $\mathbf{v}_{-i}^{(t)}$. We proceed to the proof of Lemma 4.7.
@@
For all $i\in[n]$ and $\mathbf{b}_{-i}^{(t)}\in B^{n-1}$, we denote $\delta_{i,K+1}^{(t)}(\mathbf{b}_{-i}^{(t)}):=x_{i}^{(t)}\big(1,\mathbf{b}_{-i}^{(t)}\big)$ and
@@
$\frac{\partial q_{i}^{(t)}}{\partial\pi_{k}^{(i,t)}}=\sum_{j=k}^{K+1}\delta_{i,j}^{(t)}(\mathbf{b}_{-i}^{(t)})-p^{(t)}_{i}\big(\tfrac{k-1}{K},\mathbf{b}_{-i}^{(t)}\big),\,\,\forall\,k\in[K+1].$ (17)
@@
$=$ $\Big(\sum_{j=1}^{K+1}\Big(1-\sum_{k=1}^{j}\pi_{k}^{(i,t)}\Big)\cdot\delta_{i,j}^{(t)}(\mathbf{b}_{-i}^{(t)})\Big)+\sum_{k=1}^{K+1}\pi_{k}^{(i,t)}\cdot p^{(t)}_{i}\big(\tfrac{k-1}{K},\mathbf{b}_{-i}^{(t)}\big)\qquad\qquad\,\text{(By exchanging the sums)}$
@@
$\textnormal{Rev}_{\mathcal{M}^{(t)}}\big(\mathcal{D}\mid Q^{(t)}\big)$ $=\sum_{i\in[n]}\operatorname*{\mathbb{E}}_{\mathbf{v}^{(t)}\sim\mathcal{D}}\big[p^{(t)}_{i}\big(s_{1}^{(t)}(v_{1}^{(t)}),\dots,s_{n}^{(t)}(v_{n}^{(t)})\big)\mid Q^{(t)}\big]$
@@ 4.3 Proof of the main theorem
Notice that, by Lemma 3.6, for each $i\in[n]$, the reward vectors $\nabla q_{i}^{(t)}(\bm{\pi}^{(i,t)})$ provided to bidder $i$’s input algorithm $\mathcal{A}_{i}$ are in the range $1,1]^{K+1}$, which is consistent with the standard online learning setting (Definition 4.1). Hence, by our assumption in the lemma statement, for each $i\in[n]$, algorithm $\mathcal{A}_{i}$ guarantees that
@@
5 Multiplicative weights update strikes back
In this section, we show that the multiplicative weights update (MWU) algorithm––a widely used algorithm that is manipulable (Braverman et al., 2018) and motivated the study of strategically robust online learning––simultaneously achieves the optimal regret guarantee and the best-known strategic robustness guarantee, when combined [^9] For readers familiar with online convex optimization, this combination is equivalent to applying the exponentiated gradient algorithm (Hazan, 2016, Algorithm 15) to our quantile utility functions. with our meta-algorithm (Algorithm 1).
@@ 5 Multiplicative weights update strikes back
For comparison, the standard regret guarantee of the agile and lazy online gradient descent (OGD) algorithms (Hazan, 2016, Section 5.4.1) is $O(\sqrt{T\cdot K})$. Consequently, combining the agile or lazy OGD algorithm with Algorithm 1 results in regret and strategic‐robustness guarantees with $\sqrt{T\cdot K}$ dependence. This matches the regret guarantee established by Kumar et al. (2024) for agile OGD in first-price auctions, and slightly improves their strategic robustness guarantee for multiple bidders, which scales linearly with $K$.
@@ 6 Discussion
We have shown that the quantile strategy space can be used to transform any no-regret learning algorithm into a strategically robust one, for repeated auctions satisfying allocation monotonicity and voluntary participation. It is, however, still unclear whether agile OGD itself is also strategically robust in repeated first-price auctions without relying on the quantile strategy space. Specifically, suppose all bidders run a separate instance of agile OGD for each possible value type [^10] Although the value range is a continuous interval $[0,1]$, we can discretize it into sufficiently small sub-intervals and select a representative value type from each. and use that instance to determine bids for that value type in repeated first-price auctions. Can the auctioneer strategically choose reserve prices to obtain a total revenue of $\textnormal{Mye}(\mathcal{D})\cdot T+\Omega(T)$ when facing this algorithm? We did not resolve this question in this paper. However, we note that agile OGD can incur high swap regret (Section B), and therefore, it is not strategically robust in repeated normal-form games (Mansour et al., 2022, Theorem 3).
@@ Appendix A Supplementary proofs
In this case, both bids $s^{*}(v_{i})$ and $s^{*}(v_{i}^{\prime})$ induce the same cumulative allocation. Moreover, since $\textnormal{Pair}_{i}\neq\textnormal{Pair}_{i}^{\prime}$, we must have that $\sum_{t\in[T]}p^{(t)}_{i}\big(s^{*}(v_{i}),\mathbf{b}_{-i}^{(t)}\big)\neq\sum_{t\in[T]}p^{(t)}_{i}\big(s^{*}(v_{i}^{\prime}),\mathbf{b}_{-i}^{(t)}\big)$. This means that one of the two bids, $s^{*}(v_{i})$ or $s^{*}(v_{i}^{\prime})$, incurs a strictly higher cumulative payment, even though both induce the same cumulative allocation. Hence, the bid that incurs the higher payment is strictly suboptimal for both value types $v_{i}$ and $v_{i}^{\prime}$, which contradicts our choice of $s^{*}(v_{i})$ and $s^{*}(v_{i}^{\prime})$.
@@
In this case, we have that $\sum_{t\in[T]}x_{i}^{(t)}\big(s^{*}(v_{i}),\mathbf{b}_{-i}^{(t)}\big)<\sum_{t\in[T]}x_{i}^{(t)}\big(s^{*}(v_{i}^{\prime}),\mathbf{b}_{-i}^{(t)}\big)$, since we assume that $s^{*}(v_{i})<s^{*}(v_{i}^{\prime})$, and that each $x_{i}^{(t)}$ is non-decreasing (Assumption 2.4-i). Moreover, it follows from the definition of $s^{*}(v_{i})$ that
@@
A.3 Proof of Claim 4.6
@@
Since bidders’ values $\mathbf{v}^{(t)}$ at round $t$ is sampled from $\mathcal{D}$ independently of $\tilde{\mathcal{M}}^{(t)}$, we can write $\textnormal{Rev}_{\tilde{\mathcal{M}}^{(t)}}(\mathcal{D})=\operatorname*{\mathbb{E}}_{\mathbf{v}^{(t)}\sim\mathcal{D}}\big[\sum_{i\in[n]}\tilde{p}_{i}^{(t)}(\mathbf{v}^{(t)})\big]$. We express $\tilde{p}_{i}^{(t)}(\mathbf{v}^{(t)})$ as follows,
$\tilde{p}_{i}^{(t)}(\mathbf{v}^{(t)})$ $=\tilde{x}_{i}^{(t)}\big(v_{i}^{(t)},\mathbf{v}_{-i}^{(t)}\big)v_{i}-\int_{0}^{v_{i}^{(t)}}\tilde{x}_{i}^{(t)}\big(z,\mathbf{v}_{-i}^{(t)}\big)dz$ (By Eq. (13))
@@
$=$ $\sum_{j\in[k-1]}\Big(x_{i}^{(t)}\Big(\frac{j-1}{K},\mathbf{b}_{-i}^{(t)}\Big)-x_{i}^{(t)}\Big(\frac{j}{K},\mathbf{b}_{-i}^{(t)}\Big)\Big)\cdot F_{i}^{-1}(\Pi_{j}^{(i,t)})+x_{i}^{(t)}\Big(\frac{k-1}{K},\mathbf{b}_{-i}^{(t)}\Big)v_{i}^{(t)}.$
@@
Suppose that there is only one bidder in the repeated auction game, and the bidder’s value distribution $\mathcal{D}_{1}$ is the uniform distribution over $\big[1-\frac{1}{T},1\big]$. Given any algorithm $\mathcal{A}$ that guarantees regret $\textnormal{Reg}_{\mathcal{A}}$ for the bidder in the repeated auction game, we show how to convert $\mathcal{A}$ into an algorithm $\mathcal{A}^{\prime}$ that guarantees regret $\textnormal{Reg}_{\mathcal{A}}+O(1)$ in the standard online learning setting (Definition 4.1), but with reward vectors in $\{0,1\}^{K}$. This will establish the proposition, since the worst-case regret of any algorithm in the standard online learning setting (with reward vectors in $\{0,1\}^{K}$) is at least $\Omega(\sqrt{T\log(K)})$ (Cesa-Bianchi and Lugosi, 2006, Theorem 3.7).

### v2

@@ 4.2 Main lemma
We introduce some notation to simplify the lemma statement. We let $\mathbf{e}:=(1,0,\dots,0)$ denote the first standard basis vector in $\mathbb{R}^{K+1}$. For each round $t\in[T]$, we let $Q^{(t)}:=(\bm{\pi}^{(1,t)},\dots,\bm{\pi}^{(n,t)})$ be the collection of bidders’ quantile strategies, which determine their bidding strategies $s_{i}^{(t)}$, and we let $\textnormal{Rev}_{\mathcal{M}^{(t)}}\big(\mathcal{D}\mid Q^{(t)}\big)$ denote the auctioneer’s expected revenue at round $t$ conditioned on $Q^{(t)}$, i.e.,
@@
Now we state the main lemma (Lemma 4.7), which relates the gradient of the quantile utility function to the difference between the auctioneer’s expected revenue at round $t$ and the expected revenue of the auxiliary auction. A similar relation for the special case of single-bidder first-price auctions was established by Kumar et al. (2024, page 21).
@@
We clarify that in Algorithm 1, bidder $i$’s quantile utility function $q_{i}^{(t)}$ at round $t$ (as defined in Eq. (7)) depends on the other bidders’ bids $\mathbf{b}_{-i}^{(t)}$, which in turn are determined by the other bidders’ quantile/bidding strategies and private values $\mathbf{v}_{-i}^{(t)}$ at round $t$. Hence, the expectation in Eq. (13) is taken over the randomness of $\mathbf{v}_{-i}^{(t)}$. We proceed to the proof of Lemma 4.7.
@@
For all $i\in[n]$ and $\mathbf{b}_{-i}^{(t)}\in B^{n-1}$, we denote
@@
$\frac{\partial q_{i}^{(t)}}{\partial\pi_{k}^{(i,t)}}=\Big(\sum_{j=k}^{K}\delta_{i,j}^{(t)}(\mathbf{b}_{-i}^{(t)})\Big)+p^{(t)}_{i}\big(1,\mathbf{b}_{-i}^{(t)}\big)-p^{(t)}_{i}\big(\tfrac{k-1}{K},\mathbf{b}_{-i}^{(t)}\big),\,\,\forall\,k\in[K].$ (15)
@@
$=$ $\Big(\sum_{j=1}^{K}\delta_{i,j}^{(t)}(\mathbf{b}_{-i}^{(t)})-p^{(t)}_{i}\big(0,\mathbf{b}_{-i}^{(t)}\big)\Big)-\sum_{k=1}^{K}\pi_{k}^{(i,t)}\cdot\Big(\sum_{j=k}^{K}\delta_{i,j}^{(t)}(\mathbf{b}_{-i}^{(t)})-p^{(t)}_{i}\big(\tfrac{k-1}{K},\mathbf{b}_{-i}^{(t)}\big)\Big)$
@@
$=$ $\Big(\sum_{j=1}^{K}\delta_{i,j}^{(t)}(\mathbf{b}_{-i}^{(t)})\Big)-\sum_{k=1}^{K}\pi_{k}^{(i,t)}\cdot\Big(\sum_{j=k}^{K}\delta_{i,j}^{(t)}(\mathbf{b}_{-i}^{(t)})-p^{(t)}_{i}\big(\tfrac{k-1}{K},\mathbf{b}_{-i}^{(t)}\big)\Big)$
[…]
$=$ $\Big(\sum_{j=1}^{K}\Big(1-\sum_{k=1}^{j}\pi_{k}^{(i,t)}\Big)\cdot\delta_{i,j}^{(t)}(\mathbf{b}_{-i}^{(t)})\Big)+\sum_{k=1}^{K+1}\pi_{k}^{(i,t)}\cdot p^{(t)}_{i}\big(\tfrac{k-1}{K},\mathbf{b}_{-i}^{(t)}\big)\qquad\quad\,\,\,\,\text{(By exchanging the sums)}$
@@
$\textnormal{Rev}_{\mathcal{M}^{(t)}}\big(\mathcal{D}\mid Q^{(t)}\big)$ $=\sum_{i\in[n]}\operatorname*{\mathbb{E}}_{\mathbf{v}^{(t)}\sim\mathcal{D}}\big[p^{(t)}_{i}\big(s_{1}^{(t)}(v_{1}^{(t)}),\dots,s_{n}^{(t)}(v_{n}^{(t)})\big)\mid Q^{(t)}\big]$ (By Eq. (12))
@@ 4.3 Proof of the main theorem
Notice that, by Lemma 3.6, for each $i\in[n]$, the reward vectors $\nabla q_{i}^{(t)}(\bm{\pi}^{(i,t)})$ provided to bidder $i$’s input algorithm $\mathcal{A}_{i}$ are in the range $2,1]^{K+1}$, which is consistent with the standard online learning setting (Definition 4.1). Hence, by our assumption in the lemma statement, for each $i\in[n]$, algorithm $\mathcal{A}_{i}$ guarantees that
@@
5 Instantiating the meta-algorithm with MWU
In this section, we analyze the behavior of the MWU algorithm when instantiated within our meta-algorithm (Algorithm 1). While MWU is known to be manipulable (Braverman et al., 2018), we show that it achieves the optimal regret and the best-known strategic robustness guarantee in our auction setting, when combined [^9] For readers familiar with online convex optimization, this combination is equivalent to applying the exponentiated gradient algorithm (Hazan, 2016, Algorithm 15) to our quantile utility functions. with our meta-algorithm (Algorithm 1). Importantly, the strategic robustness guarantee we obtain does not stem from any new property of MWU itself, but from the structural transformation induced by our meta-algorithm.
@@ 5 Instantiating the meta-algorithm with MWU
For comparison, the standard regret guarantee of the agile and lazy online gradient descent (OGD) algorithms (Hazan, 2016, Section 5.4.1) is $O(\sqrt{T\cdot K})$. Consequently, combining the agile or lazy OGD algorithm with Algorithm 1 results in regret and strategic‐robustness guarantees with $\sqrt{T\cdot K}$ dependence. This matches the regret guarantee established by Kumar et al. (2024) for agile OGD in first-price auctions, and slightly improves their strategic robustness guarantee for multiple bidders, which scales linearly with $K$. We note that for multiple bidders, Kumar et al. (2024) analyze a variant of agile OGD designed for unknown prior distributions, and their analysis establishes a stronger incentive compatibility property under a technical assumption.
@@ 6 Discussion
We have shown that the quantile strategy space can be used to transform any no-regret learning algorithm into a strategically robust one, for repeated auctions satisfying allocation monotonicity and voluntary participation. It is, however, still unclear whether agile OGD itself is also strategically robust in repeated first-price auctions without relying on the quantile strategy space. Specifically, suppose all bidders run a separate instance of agile OGD for each possible value type [^10] Although the value range is a continuous interval $[0,1]$, we can discretize it into sufficiently small sub-intervals and select a representative value type from each. and use that instance to determine bids for that value type in repeated first-price auctions. Can the auctioneer strategically choose reserve prices to obtain a total revenue of $\textnormal{Mye}(\mathcal{D})\cdot T+\Omega(T)$ when facing this algorithm? We did not resolve this question in this paper. However, we note that agile OGD can incur high swap regret (Section B), and therefore, it is manipulable in repeated normal-form games (Mansour et al., 2022, Theorem 3).
@@ Appendix A Supplementary proofs
In this case, both bids $s^{*}(v_{i})$ and $s^{*}(v_{i}^{\prime})$ induce the same cumulative allocation. Moreover, since $\textnormal{Pair}_{i}\neq\textnormal{Pair}_{i}^{\prime}$, we must have that
$\sum_{t\in[T]}p^{(t)}_{i}\big(s^{*}(v_{i}),\mathbf{b}_{-i}^{(t)}\big)\neq\sum_{t\in[T]}p^{(t)}_{i}\big(s^{*}(v_{i}^{\prime}),\mathbf{b}_{-i}^{(t)}\big).$
This means that one of the two bids, $s^{*}(v_{i})$ or $s^{*}(v_{i}^{\prime})$, incurs a strictly higher cumulative payment, even though both induce the same cumulative allocation. Hence, the bid that incurs the higher payment is strictly suboptimal for both value types $v_{i}$ and $v_{i}^{\prime}$, which contradicts our choice of $s^{*}(v_{i})$ and $s^{*}(v_{i}^{\prime})$.
@@
In this case, we have that
$\sum_{t\in[T]}x_{i}^{(t)}\big(s^{*}(v_{i}),\mathbf{b}_{-i}^{(t)}\big)<\sum_{t\in[T]}x_{i}^{(t)}\big(s^{*}(v_{i}^{\prime}),\mathbf{b}_{-i}^{(t)}\big),$
since we assume that $s^{*}(v_{i})<s^{*}(v_{i}^{\prime})$, and that each $x_{i}^{(t)}$ is non-decreasing (Assumption 2.4-i). Moreover, it follows from the definition of $s^{*}(v_{i})$ that
@@
A.3 Proof of Lemma 3.6
Proof of Lemma 3.6.
We first expand the function $q_{i}^{(t)}(\bm{\pi})$ as follows,
$q_{i}^{(t)}(\bm{\pi})$ $=u^{(t)}_{i}(s_{i,\bm{\pi}})$ (By Eq. (7))
$=\sum_{j\in[K]}\int_{F_{i}^{-1}(\sum_{\ell=1}^{j-1}\pi_{\ell})}^{F_{i}^{-1}(\sum_{\ell=1}^{j}\pi_{\ell})}x_{i}^{(t)}\big(\tfrac{j-1}{K},\mathbf{b}_{-i}^{(t)}\big)\cdot v_{i}-p^{(t)}_{i}\big(\tfrac{j-1}{K},\mathbf{b}_{-i}^{(t)}\big)dF_{i}(v_{i})$
$\quad+\int_{F_{i}^{-1}(\sum_{\ell=1}^{K}\pi_{\ell})}^{1}x_{i}^{(t)}\big(1,\mathbf{b}_{-i}^{(t)}\big)\cdot v_{i}-p^{(t)}_{i}\big(1,\mathbf{b}_{-i}^{(t)}\big)dF_{i}(v_{i})$ (By Eq. (6))
$=\sum_{j\in[K]}\int_{\sum_{\ell=1}^{j-1}\pi_{\ell}}^{\sum_{\ell=1}^{j}\pi_{\ell}}x_{i}^{(t)}\big(\tfrac{j-1}{K},\mathbf{b}_{-i}^{(t)}\big)\cdot F_{i}^{-1}(z)-p^{(t)}_{i}\big(\tfrac{j-1}{K},\mathbf{b}_{-i}^{(t)}\big)dz$
$\quad+\int_{\sum_{\ell=1}^{K}\pi_{\ell}}^{1}x_{i}^{(t)}\big(1,\mathbf{b}_{-i}^{(t)}\big)\cdot F_{i}^{-1}(z)-p^{(t)}_{i}\big(1,\mathbf{b}_{-i}^{(t)}\big)dz.$ (23)
Then, we define the partial-sum functions $\tau_{j}(\bm{\pi}):=\sum_{\ell=1}^{j}\pi_{\ell}$ for all $j\in[K]$, and we define an intermediate function $\sigma_{i}^{(t)}:[0,1]^{K}\to[-1,1]$ as follows [^11] We note that $\sigma_{i}^{(t)}$ is essentially a generalization of bidders’ utility functions in Kumar et al. (2024, Eq. (3)) from first-price auctions to arbitrary auctions.,
$\sigma_{i}^{(t)}(y_{1},\dots,y_{K}):=\sum_{j\in[K+1]}\int_{y_{j-1}}^{y_{j}}x_{i}^{(t)}\big(\tfrac{j-1}{K},\mathbf{b}_{-i}^{(t)}\big)\cdot F_{i}^{-1}(z)-p^{(t)}_{i}\big(\tfrac{j-1}{K},\mathbf{b}_{-i}^{(t)}\big)dz,$
where we let $y_{0}:=0$ and $y_{K+1}:=1$. We notice that $q_{i}^{(t)}(\bm{\pi})=\sigma_{i}^{(t)}(\tau_{1}(\bm{\pi}),\dots,\tau_{K}(\bm{\pi}))$ by Eq. (A.3), and moreover, for each $j\in[K]$, $\tau_{j}$ is a linear function of $\bm{\pi}$.
Concavity of $q_{i}^{(t)}$
Since linear transformations preserve concavity (Boyd and Vandenberghe, 2004, Section 3.2), to show that $q_{i}^{(t)}$ is concave, it suffices to prove that $\sigma_{i}^{(t)}$ is concave. To this end, we observe that the partial derivatives of $\sigma_{i}^{(t)}$ with respect to $y_{j}$, for all $j\in[K]$, are given by
$\frac{\partial\sigma_{i}^{(t)}}{\partial y_{j}}$ $=\big(x_{i}^{(t)}\big(\tfrac{j-1}{K},\mathbf{b}_{-i}^{(t)}\big)-x_{i}^{(t)}\big(\tfrac{j}{K},\mathbf{b}_{-i}^{(t)}\big)\big)\cdot F_{i}^{-1}(y_{j})-\big(p^{(t)}_{i}\big(\tfrac{j-1}{K},\mathbf{b}_{-i}^{(t)}\big)-p^{(t)}_{i}\big(\tfrac{j}{K},\mathbf{b}_{-i}^{(t)}\big)\big),\,\,\forall\,j\in[K].$ (24)
We notice that for all $j\in[K]$, the partial derivative $\frac{\partial\sigma_{i}^{(t)}}{\partial y_{j}}$ depends only on $y_{j}$. Thus, to prove that $\sigma_{i}^{(t)}$ is concave, it suffices to show that for all $j\in[K]$, $\frac{\partial\sigma_{i}^{(t)}}{\partial y_{j}}$ is a non-increasing function of $y_{j}$. We observe that $F_{i}^{-1}(y_{j})$ is non-decreasing in $y_{j}$ since $F_{i}$ is a CDF, and that $x_{i}^{(t)}\big(\tfrac{j-1}{K},\mathbf{b}_{-i}^{(t)}\big)-x_{i}^{(t)}\big(\tfrac{j}{K},\mathbf{b}_{-i}^{(t)}\big)$ is non-positive because $x_{i}^{(t)}$ is non-decreasing by Assumption 2.4-i. It follows that $\frac{\partial\sigma_{i}^{(t)}}{\partial y_{j}}$ is a non-increasing function of $y_{j}$ for all $j\in[K]$, which implies that $\sigma_{i}^{(t)}$, and hence $q_{i}^{(t)}$, is concave.
Gradient of $q_{i}^{(t)}$
Next, we note that $\frac{\partial q_{i}^{(t)}}{\partial\pi_{K+1}}=0$ since $q_{i}^{(t)}$ is independent of $\pi_{K+1}$. Moreover, since $q_{i}^{(t)}(\bm{\pi})=\sigma_{i}^{(t)}(\tau_{1}(\bm{\pi}),\dots,\tau_{K}(\bm{\pi}))$, for each $k\in[K]$, we can compute the partial derivative of $q_{i}^{(t)}$ with respect to $\pi_{k}$ as follows,
$\frac{\partial q_{i}^{(t)}}{\partial\pi_{k}}$ $=\sum_{j=1}^{K}\frac{\partial\sigma_{i}^{(t)}}{\partial\tau_{j}}\cdot\frac{\partial\tau_{j}}{\partial\pi_{k}}=\sum_{j=k}^{K}\frac{\partial\sigma_{i}^{(t)}}{\partial\tau_{j}}$
(By Eq. (24))
$=\bigg(\sum_{j=k}^{K}\big(x_{i}^{(t)}\big(\tfrac{j-1}{K},\mathbf{b}_{-i}^{(t)}\big)-x_{i}^{(t)}\big(\tfrac{j}{K},\mathbf{b}_{-i}^{(t)}\big)\big)\cdot F_{i}^{-1}(\tau_{j}(\bm{\pi}))\bigg)+p_{i}^{(t)}\big(1,\mathbf{b}_{-i}^{(t)}\big)-p^{(t)}_{i}\big(\tfrac{k-1}{K},\mathbf{b}_{-i}^{(t)}\big)$
$\qquad\text{(By a telescoping sum)},$
which establishes Eq. (9). To bound $\frac{\partial q_{i}^{(t)}}{\partial\pi_{k}}$ for any $k\in[K]$, we notice that each term $x_{i}^{(t)}\big(\tfrac{j-1}{K},\mathbf{b}_{-i}^{(t)}\big)-x_{i}^{(t)}\big(\tfrac{j}{K},\mathbf{b}_{-i}^{(t)}\big)$ is non-positive (by Assumption 2.4-i) and $F_{i}^{-1}(\tau_{j}(\bm{\pi}))\in[0,1]$. Hence, we have that
Combining this with Eq. (9), we obtain that
$\frac{\partial q_{i}^{(t)}}{\partial\pi_{k}}$ $\leq p_{i}^{(t)}\big(1,\mathbf{b}_{-i}^{(t)}\big)-p^{(t)}_{i}\big(\tfrac{k-1}{K},\mathbf{b}_{-i}^{(t)}\big)\leq 1,$
$\frac{\partial q_{i}^{(t)}}{\partial\pi_{k}}$ $\geq x_{i}^{(t)}\big(\tfrac{k-1}{K},\mathbf{b}_{-i}^{(t)}\big)-x_{i}^{(t)}\big(1,\mathbf{b}_{-i}^{(t)}\big)+p_{i}^{(t)}\big(1,\mathbf{b}_{-i}^{(t)}\big)-p^{(t)}_{i}\big(\tfrac{k-1}{K},\mathbf{b}_{-i}^{(t)}\big)$
$\geq-2,$
which completes the proof of the lemma. ∎
A.4 Proof of Claim 4.6
@@
Since bidders’ value vector $\mathbf{v}^{(t)}$ at round $t$ is sampled from $\mathcal{D}$ independently of $\tilde{\mathcal{M}}^{(t)}$, we can write $\textnormal{Rev}_{\tilde{\mathcal{M}}^{(t)}}(\mathcal{D})=\operatorname*{\mathbb{E}}_{\mathbf{v}^{(t)}\sim\mathcal{D}}\big[\sum_{i\in[n]}\tilde{p}_{i}^{(t)}(\mathbf{v}^{(t)})\big]$. We express $\tilde{p}_{i}^{(t)}(\mathbf{v}^{(t)})$ as follows,
$\tilde{p}_{i}^{(t)}(\mathbf{v}^{(t)})$ $=\tilde{x}_{i}^{(t)}\big(v_{i}^{(t)},\mathbf{v}_{-i}^{(t)}\big)v_{i}^{(t)}-\int_{0}^{v_{i}^{(t)}}\tilde{x}_{i}^{(t)}\big(z,\mathbf{v}_{-i}^{(t)}\big)dz$ (By Eq. (11))
@@
$=$ $\sum_{j\in[k-1]}\Big(x_{i}^{(t)}\Big(\frac{j-1}{K},\mathbf{b}_{-i}^{(t)}\Big)-x_{i}^{(t)}\Big(\frac{j}{K},\mathbf{b}_{-i}^{(t)}\Big)\Big)\cdot F_{i}^{-1}(\Pi_{j}^{(i,t)})+x_{i}^{(t)}\Big(\frac{k-1}{K},\mathbf{b}_{-i}^{(t)}\Big)v_{i}^{(t)}$
$\quad\text{(By $F_{i}^{-1}(\Pi_{0}^{(i,t)})=F_{i}^{-1}(0)=0$ and rearranging)}.$
@@
Suppose that there is only one bidder in the repeated auction game, and the bidder’s value distribution $\mathcal{D}_{1}$ is the uniform distribution over $\big[1-\frac{1}{T},1\big]$ [^12] The proof applies to any continuous distribution with sufficiently large probability mass on $\big[1-\frac{1}{T},1\big]$.. Given any algorithm $\mathcal{A}$ that guarantees regret $\textnormal{Reg}_{\mathcal{A}}$ for the bidder in the repeated auction game, we show how to convert $\mathcal{A}$ into an algorithm $\mathcal{A}^{\prime}$ that guarantees regret $\textnormal{Reg}_{\mathcal{A}}+O(1)$ in the standard online learning setting (Definition 4.1), but with reward vectors in $\{0,1\}^{K}$. This will establish the proposition, since the worst-case regret of any algorithm in the standard online learning setting (with reward vectors in $\{0,1\}^{K}$) is at least $\Omega(\sqrt{T\log(K)})$ (Cesa-Bianchi and Lugosi, 2006, Theorem 3.7).